Each breakdown of a graphic plotter table at Airbus Industries costs

$ 50$50.

Following is the historical breakdown​ distribution:

                                                                                                            

Number of breakdowns

0

1

2

3

4

Daily breakdown probability

0.120.12

0.210.21

0.390.39

0.190.19

0.090.09

For Airbus​ Industries, the expected daily breakdown cost​ =

nothing

dollars per day ​(round your response to two decimal​ places).
Answer #1

Expected number of breakdown = sum of ( Number of breakdowns* Daily breakdown probability)

= 0*0.12+ 1*0.21+ 2*0.39+3*0.19+ 4*.09

=1.92

Cost of one breakdown $50

So cost of breakdown - expected breakdown * cost of one breakdown= 50*1.92 = $96